OpenStudy (anonymous):

In triangle ABC, centroid D is on median (of line) AM. Ad = x + 5 and DM = 3x - 5. Find AM.

OpenStudy (tad1):

x+5 = 3x - 5 10 = 2x x = 5 so lengtf of AM = 10 + 10 = 20

OpenStudy (anonymous):

That's wrong, the only possible answers are 12, 6, 13, and 4.